# Transfer Learning for High-dimensional Quantile Regression with Distribution Shift

### v1

@@ 5.1 A Practical Guide for Tuning Parameters
To select the $\ell_{1}$-QR tuning parameters, we deviate from the vanilla penalized quantile regression and instead follow the $\lambda$-construction idea in Belloni and Chernozhukov, (2011); Giessing and Wang, (2023). Specifically, the regularization term changes from $\lambda\|\boldsymbol{\beta}\|_{1}$ to $\lambda\sqrt{\tau(1-\tau)}\sum_{j=1}^{p}\widehat{\sigma}_{k,j}|\beta_{j}|$, with $\widehat{\sigma}_{k,j}^{2}:=n_{k}^{-1}\sum_{i=1}^{n_{k}}(\boldsymbol{x}_{ij}^{(k)})^{2}$ and $\lambda=c^{*}\cdot\Lambda_{\tau}^{(k)}(\alpha^{*}|\{\boldsymbol{x}_{i}^{(k)}\}_{i=1}^{n_{k}})$, where $\Lambda_{\tau}^{(k)}(\alpha^{*}|\{\boldsymbol{x}_{i}^{(k)}\}_{i=1}^{n_{k}})$ is the $(1-\alpha^{*})$-th quantile of $\Lambda_{\tau}^{(k)}|\{\boldsymbol{x}_{i}^{(k)}\}_{i=1}^{n_{k}}$ and
@@
with $U_{1},\ldots,U_{n_{k}}$ are i.i.d. Uniform$(0,1)$ random variables, independent of the covariates $\{\boldsymbol{x}_{i}^{(k)}\}_{i=1}^{n_{k}}$. Such a pilot construction can save the computation cost a lot compared to the cross-validation. $\alpha^{*}$ is set as 0.05 uniformly, while we set $c^{*}=1$ for $\lambda_{0},\lambda_{\boldsymbol{\beta^{*}}}$, $c^{*}=1.5$ for $\{\lambda_{k}\}_{k=1}^{K}$, $c^{*}=1$ for $\{\lambda_{0.5}^{(k)}\}_{k\in\{0\}\cup\widehat{\mathcal{C}}}$, $c^{*}=0.95$ for $\{\lambda_{0.7}^{(k)}\}_{k\in\{0\}\cup\widehat{\mathcal{C}}}$ and $c^{*}=0.9$ for $\{\lambda_{0.2}^{(k)}\}_{k\in\{0\}\cup\widehat{\mathcal{C}}}$. This procedure can be executed through R functions LassoLambdaHat() in the quantreg package.
For the bandwidth parameters $\{b_{k}\}_{k=0}^{K}$, we follow the construction details in Huang et al., (2023) with estimators $\widetilde{\boldsymbol{\beta}},\{\widetilde{\boldsymbol{\delta}}^{(k)}\}_{k=1}^{K}$ in Algorithm 2. We set the threshold parameter $t_{1}=5,t_{2}=0.3$ in Algorithm 2 for simplicity, although one can still choose $t_{1},t_{2}$ by cross-validation. In our simulations the performance of transfer estimators is not sensitive to the choice of $t_{1},t_{2}$ as long as they fall into a proper interval. We also tried other values and the results are similar. For simplicity, we set the post-selected threshold $\bar{\lambda}=0.01$. Then for the tuning parameter $\lambda_{\boldsymbol{\theta}}^{(k)}$ in the projection, we use the common high-dimensional choice $c_{\boldsymbol{\theta}}\sqrt{\log p/n_{k}}$ in the experiments, where the constant $c_{\boldsymbol{\theta}}$ is set as $0.1$.
5.2 Estimation Errors
We follow a similar setup as in Zhang and Zhu, (2023); Huang et al., (2023). Consider a target study with size $n_{0}=200$, dimension $p=500$, and sparsity level $s=10$. We simulate two types of target model:
[…]
where $\boldsymbol{\beta^{*}}=(\mathbf{1}_{s},\mathbf{0}_{p-s})^{\top}$, $\boldsymbol{x}_{i}^{(0)}\sim\mathcal{N}(\mathbf{0}_{p},\boldsymbol{\Sigma}_{\boldsymbol{x}})$ with $\boldsymbol{\Sigma}_{\boldsymbol{x}}=\left(0.7^{|i-j|}\right)_{1\leq i,j\leq p}$, and $\epsilon_{i}^{(0)}$ follows the shifted standard normal distribution with quantile levels $\tau\in\{0.2,0.5,0.7\}$.
The source study number $K=5$ and follows the homogeneous/heterogeneous model type (16) with $(y_{i}^{(k)},\boldsymbol{x}_{i}^{(k)},\boldsymbol{w}^{(k)},\epsilon_{i}^{(k)})_{i=1}^{n_{k}}$. To simulate the distribution shift, we consider the following parameter configurations for the source studies.
  • Covariate shift. For each source study $k=1,\ldots,K$, we set each covariate $\boldsymbol{x}_{i}^{(k)}\sim\mathcal{N}\left(\mathbf{0}_{p},\boldsymbol{\Sigma}_{\boldsymbol{x}}+\boldsymbol{\epsilon}_{\boldsymbol{x}}\boldsymbol{\epsilon}_{\boldsymbol{x}}^{T}\right)$ with $\boldsymbol{\epsilon}_{\boldsymbol{x}}\sim\mathcal{N}\left(\mathbf{0}_{p},0.3^{2}\boldsymbol{I}_{p}\right)$.
  • Parameter shift. For each source study $k=1,\ldots,K$, denote $\zeta_{i}^{(k)}$ as independent Rademacher random variables (taking values in $\{1,-1\}$ with equal probability). Following the notation in Section 3, we denote the first oracle set $\mathcal{C}_{h_{1}}$, in which the $\ell_{1}$-norm of parameter contrast $\|\boldsymbol{\delta}^{(k)}\|_{1}=\|\boldsymbol{w}^{(k)}-\boldsymbol{\beta^{*}}\|_{1}$ of each source study is relatively small. We set
    $w_{j}^{(k)}=\begin{cases}\beta^{*}_{j}+h_{1}/100\cdot r_{j}^{(k)}\cdot\mathbb{I}\left\{j\in H^{(k)}\cup[s/2]\right\},&k\in\mathcal{C}_{h_{1}},\\ h_{1}/10\cdot r_{j}^{(k)}\cdot\mathbb{I}\left\{j\in H^{(k)}\cup[s/2]\right\},&k\in\mathcal{C}_{h_{1}}^{c},\end{cases}$
    where $[s/2]=\{1,\ldots,s/2\}$, $H^{(k)}$ is a random subset chosen from $\{s/2+1,\ldots,p\}$ with $|H^{(k)}|=50$. It can be verified that $\max_{k\in\mathcal{C}_{h_{1}}}\|\boldsymbol{\delta}^{(k)}\|_{1}<h_{1}<\min_{k\in\mathcal{C}_{h_{1}}^{c}}\|\boldsymbol{\delta}^{(k)}\|_{1}$. In this setting, we consider $h_{1}=5$.
  • Residual shift. For each source study $k=1,\ldots,K$, the residual $r_{i}^{(k)}$ either follows the standard normal distribution, or it follows one type of the following abnormal distributions:
    1. 1.
      (Cauchy) cauchy distribution: $\mathcal{C}(0,3)$;
    2. 2.
      (Mixed) mixed gaussian: $z\mathcal{N}(-3,0.5)+(1-z)\mathcal{N}(3,0.5)$, where $z\sim\mathcal{B}ernoulli(\tau)$;
    3. 3.
      (Noisy) noisy setting: $\mathcal{N}(0,5^{2})$.
    All distributions have already been shifted to meet $\mathbb{P}(\epsilon_{i}^{(k)}\leq 0|\boldsymbol{x}_{i}^{(k)})=\tau$. We denote the set containing all standard normal source studies by $\mathcal{C}_{h_{2}}$, where we can let $h_{2}=4$ satisfy $\max_{k\in\mathcal{C}_{h_{2}}}\mathbb{E}[f_{i}^{(0)}]/\mathbb{E}[f_{i}^{(k)}]<h_{2}<\min_{k\in\mathcal{C}_{h_{2}}^{c}}\mathbb{E}[f_{i}^{(0)}]/\mathbb{E}[f_{i}^{(k)}]$. To clearly demonstrate the negative impact of residual shift, we set the sample size to $n_{k}=100\cdot\mathbb{I}\{k\in\mathcal{C}_{h_{2}}\}+200\cdot\mathbb{I}\{k\in\mathcal{C}_{h_{2}}^{c}\}$.
Based on the definition in (3), we define $\mathcal{C}_{h}=\mathcal{C}_{h_{1}}\cap\mathcal{C}_{h_{2}}$ as the transferable oracle set. To capture both parameter and residual shift in one setting, we vary the parametric transferable set $\mathcal{C}_{h_{1}}$ from 1 to 5, and assume that each source study follows a divergent distribution with probability 1/2. For example, in the “Cauchy” setting, each $\epsilon_{i}^{(k)}$ either follows a normal distribution like $\epsilon_{i}^{(0)}$ or a Cauchy distribution with equal probability. The same applies to the “Mixed” and “Noisy” settings. We compare six specific methods, which are: 1) L1-QR: conduct $\ell_{1}$ quantile regression only on the target data; 2) Pooling: Algorithm 1 on all source studies; 3) Oracle: Algorithm 1 on the oracle transferable set $\mathcal{C}_{h}$; 4) TransQR: Algorithm 2, our transfer approach; 5) Oracle_PS: Algorithm 1 on $\mathcal{C}_{h_{1}}$ (only considering parameter shift), the baseline in the previous literature; 6) TransQR_2step: two-step transfer approach proposed in Huang et al., (2023).
We evaluate the performance of each method in terms of the $\ell_{2}$-error $\|\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta}\|_{2}^{2}$, the $\ell_{1}$-error $\|\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta}\|_{1}$ and the average prediction error $n_{0}^{-1}\sum_{i=1}^{n_{0}}[(\boldsymbol{x}_{i}^{t})^{\top}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})]^{2}$, where the test samples $\{\boldsymbol{x}_{i}^{t}\}_{i=1}^{n_{0}}$ generate in the same way of $\{\boldsymbol{x}_{i}^{(0)}\}_{i=1}^{n_{0}}$. We repeat the experiment for 100 times and average the results for each setting respectively, where the average $\ell_{2}$-errors are shown in Figure 5.1 and 5.2. More simulation results are reported in the supplementary materials.
Figure 5.1: $\ell_{2}$-estimation errors of various methods under homogeneous model at quantile levels $\tau=\{0.2,0.5,0.7\}$.
Figure 5.2: $\ell_{2}$-estimation errors of various methods under heterogeneous model at quantile levels $\tau=\{0.2,0.5,0.7\}$
From Figure 5.1 and 5.2, we highlight three points.
  1. 1.
    As we expect, the Oracle approach has the best performance across all scenarios because it transfers knowledge exactly from the ideal source studies. Moreover, more transferable source studies can produce smaller estimation errors while they do not affect the performance of the single task L1-QR. This fact corresponds our motivation that one can benefit from transfer learning, and conforms with our theoretical results.
  2. 2.
    Similar to the previous work, the Pooling approach doesn’t perform well when there are scarce transferable source studies. This phenomenon strongly points out the importance of the source detection step before knowledge transfer. It needs to emphasize that the residual shift indeed can influence the transfer performance, since the Oracle_PS method performs much worser than the Oracle method. Note that Oracle_PS is also conducted on the prespecified studies, which is actually regarded as the oracle method in all previous literature. The only difference from the Oracle is that Oracle_PS only considers parameter shift, i.e., small $\|\boldsymbol{\delta}^{(k)}\|_{1}$ for screening transferable sets, which demonstrates that our proposed $\tau$-th density ratio is indeed an important index for knowledge transfer under residual shift.
  3. 3.
    It is amazing that our proposed method TransQR almost match the oracle method, while the two-step approach TransQR_2step is not satisfactory. The failure of the two-step transfer may be due to its sensitivity to covariate shift, as well as its detection procedure’s inability to effectively screen out the source studies with residual shift. This fact reflects that our method is more effective for transferable set detection and parameter estimation under distribution shift.
5.3 Inference Results
To give some intuitive understandings of Algorithm 3, we conduct simulations for statistical inference on $\beta^{*}_{1}$ under the heterogeneous model setting in (16) with some adaptive changes. We let $|\mathcal{C}_{h_{1}}|=5,h_{1}=10$ with varying $|\delta_{1}^{(k)}|\in\{0.01,0.1\}$ (controlling approximation error $r_{i}^{(k)}$), also $n_{0}=n_{k|k\in\mathcal{C}_{h_{2}}}=300$ and $n_{k|k\in\mathcal{C}_{h_{2}}^{c}}=500$. In the simulations above we have verified the detection consistency, hence for computational simplicity we assume $\mathcal{C}_{h_{2}}:=\{1,2,3\}$ is known, and assume the set in $\mathcal{C}_{h_{2}}^{c}$ follows the mixed gaussian distribution.
We compare five specific methods, including:
  1. 1.
    Debias_tar: debiasing on the target study by Algorithm 3 with the L1-QR estimate (baseline);
  2. 2.
    Debias_trans_pool: debiasing on all studies by Algorithm 3 with the Pooling estimate;
  3. 3.
    Debias_trans_tar: debiasing on the target study by Algorithm 3 with the TransQR estimate;
  4. 4.
    Debias_trans: debiasing on the target and transferable source studies by Algorithm 3 with the TransQR estimate;
  5. 5.
    Debias_dl: debiased Lasso approach proposed in Huang et al., (2023) with the TransQR_2step estimate.
Note that Debias_trans is our primary approach for comparison. However, if the quality control step (Algorithm 4) rejects the normality test, we recommend using a more conservative approach, namely Debias_trans_tar. As will be demonstrated later, when $|\delta_{1}^{(k)}|=0.1$, Debias_trans fails due to the non-negligible approximation error $r_{i}^{(k)}$, whereas Debias_trans_tar proves to be the more robust option. We assess the inference efficiency of various methods across four dimensions: (a) coverage probability of 95% confidence intervals, (b) bias comparison, (c) average length of 95% confidence intervals, and (d) density plots of normalized estimates against $\mathcal{N}(0,1)$. The experiment was repeated 1,000 times using three methods at $\tau\in\{0.2,0.5,0.7\}$. The results (a) are summarized in Table 1-2. The results (b)-(d) are summarized in Figure 5.3-5.4.

### v2

@@ 8.1.3 A Practical Guide for Tuning Parameters
  • To select the $\ell_{1}$-QR tuning parameters, we deviate from the vanilla penalized quantile regression and instead follow the $\lambda$-construction idea in [belloni2011l1, giessing2023debiased]. Specifically, the regularization term changes from $\lambda\|\boldsymbol{\beta}\|_{1}$ to $\lambda\sqrt{\tau(1-\tau)}\sum_{j=1}^{p}\widehat{\sigma}_{k,j}|\beta_{j}|$, with $\widehat{\sigma}_{k,j}^{2}:=n_{k}^{-1}\sum_{i=1}^{n_{k}}(\boldsymbol{x}_{ij}^{(k)})^{2}$ and $\lambda=c^{*}\cdot\Lambda_{\tau}^{(k)}(\alpha^{*}|\{\boldsymbol{x}_{i}^{(k)}\}_{i=1}^{n_{k}})$, where $\Lambda_{\tau}^{(k)}(\alpha^{*}|\{\boldsymbol{x}_{i}^{(k)}\}_{i=1}^{n_{k}})$ is the $(1-\alpha^{*})$-th quantile of $\Lambda_{\tau}^{(k)}|\{\boldsymbol{x}_{i}^{(k)}\}_{i=1}^{n_{k}}$ and
@@
    with $U_{1},\ldots,U_{n_{k}}$ are i.i.d. Uniform(0,1)(0,1) random variables, independent of the covariates $\{\boldsymbol{x}_{i}^{(k)}\}_{i=1}^{n_{k}}$. Such a pilot construction can save the computation cost a lot compared to the cross-validation. $\alpha^{*}$ is set as 0.05 uniformly, while we set $c^{*}=1$ for $\lambda_{0},\lambda_{\boldsymbol{\beta^{*}}}$, $c^{*}=1.5$ for $\{\lambda_{k}\}_{k=1}^{K}$, $c^{*}=2\sqrt{\tau(1-\tau)}$ for $\{\lambda_{\tau}^{(k)}\}_{k\in\{0\}\cup\widehat{\mathcal{C}}}$. This procedure can be executed through R functions LassoLambdaHat() in the quantreg package.
  • For the bandwidth parameters $\{b_{k}\}_{k=0}^{K}$, we follow the Powell bandwidth construction [dai2021inference, huang2023estimation] with estimators $\widetilde{\boldsymbol{\beta}},\{\widetilde{\boldsymbol{\delta}}^{(k)}\}_{k=1}^{K}$ in Algorithm 1. Specifically,
    $b_{k}=\left(\Phi^{-1}(\tau+\widetilde{b}_{k})-\Phi^{-1}(\tau-\widetilde{b}_{k})\right)\min\left\{\sqrt{\widehat{var}\left[\widehat{\boldsymbol{e}}^{(k)}_{i}\right]},\frac{\widehat{Q}_{0.75}\left[\widehat{\boldsymbol{e}}^{(k)}_{i}\right]-\widehat{Q}_{0.25}\left[\widehat{\boldsymbol{e}}^{(k)}_{i}\right]}{1.34}\right\},$
    $\widetilde{b}_{k}=n_{k}^{-1/3}\left\{\Phi^{-1}(0.975)\right\}^{2/3}\left(\frac{1.5\phi^{2}(\Phi^{-1}(\tau))}{2\Phi^{-2}(\tau)+1}\right)^{1/3},$
    where $\widehat{\boldsymbol{e}}^{(k)}_{i}=y_{i}^{(k)}-(\boldsymbol{x}_{i}^{(k)})^{\top}(\widetilde{\boldsymbol{\beta}}+\widetilde{\boldsymbol{\delta}}^{(k)})$, $\widehat{var}[\widehat{\boldsymbol{e}}^{(k)}_{i}$] and $\widehat{Q}_{\tau}[\widehat{\boldsymbol{e}}^{(k)}_{i}$] denote the sample variance and the $\tau$-th sample quantile of $\{\widehat{\boldsymbol{e}}^{(k)}_{i}\}_{i=1}^{n_{k}}$.
  • For the threshold parameter $t_{1},t_{2}$ in Algorithm 1, one practical strategy is to conduct the cross-validation procedure [tian2022transfer, zhang2022transfer] for a target study and several candidate source studies, where we evaluate a grid of threshold values and choose the thresholds that yield the best prediction performance (least quantile loss on the validation set). The chosen thresholds can then be applied across all related transfer learning tasks. For moderately complex scenarios, this procedure is not computationally expensive and provides sufficiently accurate threshold choices. To further examine the impact of threshold selection, we conducted sensitivity analyses by varying the values of $t_{1}$ and $t_{2}$ in the simulations, as shown in Figures 13 and 14. The results indicate that as long as the thresholds lie within a reasonable interval, our method’s performance remains close to the oracle benchmark, consistent with the detection consistency established in Theorem 3. For all simulations in this paper, we set $t_{1}=5$ and $t_{2}=0.3$.
  • For the post-selected threshold $\bar{\lambda}$, we directly let it be 0.01 for simplicity.
  • For the parameters $\lambda_{\boldsymbol{\theta}}^{(k)}$ in the projection, we use the common pivotal choice $c_{\boldsymbol{\theta}}\sqrt{n_{k}\log p}$ in the high-dimensional settings. For different simulations (with settings described in Section 8.2), the constant $c_{\boldsymbol{\theta}}$ is set as $0.1$ for the target study and source studies under normal distributions without strong covariate shift, $0.05$ for source studies under normal distributions with strong covariate shift, $0.08$ for the target study and source studies under $t$ distributions without strong covariate shift, $0.03$ for source studies under $t$ distributions with strong covariate shift.
8.1.4 Quality Control for Statistical Inference
In Section 4, we introduce the Neyman orthogonal debiased approach to incorporate informative source studies for statistical inference. when the source studies in the oracle transferable set $\mathcal{C}_{h}$ satisfy the requirement for Theorem 4, we can get the asymptotic result of Algorithm 2 by combining Theorem 3 and 4.
Corollary 2.
Assume all conditions required by Theorem 1, 3 and 4 hold for the oracle transferable set $\mathcal{C}_{h}$. For the estimator $\widehat{\alpha}$ obtained from Algorithm 2, we have
$\sigma_{\mathcal{C}_{h}}^{-1}\sqrt{n_{\mathcal{C}_{h}}}(\widehat{\alpha}-\beta_{1}^{*})=\mathbb{U}_{\mathcal{C}_{h}}(\tau)+o_{p}(1),\hskip 18.49988pt\mathbb{U}_{\mathcal{C}_{h}}(\tau)\leadsto\mathcal{N}(0,1),$
where $\sigma_{\mathcal{C}_{h}},\mathbb{U}_{\mathcal{C}_{h}}$ follow the definitions of $\sigma_{\mathcal{C}},\mathbb{U}_{\mathcal{C}}$ in Theorem 4 with $\mathcal{C}=\mathcal{C}_{h}$. Moreover, the estimator $\widehat{\sigma}_{\widehat{\mathcal{C}}}^{2}=\tau(1-\tau)n_{\widehat{\mathcal{C}}}[\sum_{k\in\{0\}\cup\widehat{\mathcal{C}}}\sum_{i=1}^{n_{k}}(\widehat{v}_{i}^{(k)})^{2}]^{-1}$ is consistent.
Although this corollary seems to provide the theoretical guarantee for the post-detection inference procedure, leveraging source data still needs parameter shift small enough, especially for the approximation error $r_{i}^{(k)}=x_{i,1}^{(k)}\delta^{(k)}_{1},k\in\mathcal{C}_{h}$. In fact, none of the existing debiased approaches can be applied to the approximate error term, because each $\delta^{(k)}_{1}$ shares the same covariate $x_{i,1}^{(k)}$ with the target parameter $\beta^{*}_{1}$. As noted by [belloni2019valid], the orthogonal approach remains theoretically valid as long as $\mathbb{E}[(r_{i}^{(k)})^{2}]=(\mathbb{E}[x_{i1}^{(k)}]\delta_{1}^{(k)})^{2}$ stays below a certain threshold. In the transfer setting, introducing more external data increases the requirement for the contrast $\delta_{1}^{(k)}$, where Condition 7 showed that a sufficient theoretical bound is $o(n_{\widehat{\mathcal{C}}}^{-1/2})$. This requirement is understandable since we need smaller parameter contrast for better normality results. However, constructing consistent selection procedures or hypothesis tests for this condition is nearly impossible, especially when dynamically selecting $\widehat{\mathcal{C}}$. Although there is no theoretical guarantee, we propose a practical quality control procedure to validate the transfer debiased estimator. We can use prior knowledge to exclude certain source studies with non-negligible $\widetilde{\delta}_{1}^{(k)}$ and perform a bootstrap test to check the normality of the debiased estimator. If the normality hypothesis is rejected, we revert to debiasing only the target study, yielding a result similar to previous approaches. The full procedure is outlined in Algorithm 4.
Input: Same as Algorithm 2.
  1. A)
    Pre-control: for each source study $k\in\widehat{\mathcal{C}}$, if there has prior knowledge that the parameter contrast $\delta_{1}^{(k)}$ is non-negligible, or $|\widetilde{\delta}_{1}^{(k)}|$ is larger than a prespecified constant, then we remove this study. Denote the new detected set by $\widetilde{\mathcal{C}}$.
  2. B)
    Post-control: normality test with quantile bootstrap.
    1. (1)
      Conduct the quantile pair bootstrap [hahn1995bootstrapping] $R$ times on the data $\{((\boldsymbol{x}_{i}^{(k)})^{\top},y_{i}^{(k)})_{i=1}^{n_{k}}\}_{k\in\{0\}\cup\widetilde{\mathcal{C}}}$ to obtain the debiased estimates $\{\widehat{\alpha}_{r}\}_{r=1}^{R}$.
    2. (2)
      Check the normality of $\{\widehat{\alpha}_{r}\}_{r=1}^{R}$ by Shapiro-Wilk test [shapiro1965analysis].
- If rejecting null hypothesis ($p\text{-value}<0.05$), we choose to only debias on the target study by $\arg\min_{\alpha\in\mathcal{A}_{0}}|\sum_{i=1}^{n_{0}}\widehat{\psi}_{i}^{(0)}(\alpha)|$, where the search region $\mathcal{A}_{0}=\{\alpha\in\mathbb{R}:|\alpha-\widehat{\beta}_{1}|\leq 10[n_{0}^{-1}\sum_{i=1}^{n_{0}}(x^{(0)}_{i,1})^{2}]^{-1/2}/\log n_{0}\}$, which is exactly the approach in [belloni2019valid].
- If not rejecting null hypothesis, conduct Algorithm 2 with $\widetilde{\mathcal{C}}$. Output: Debiased estimate $\widehat{\alpha}$.
Algorithm 4 Quality control for Algorithm 2
In Section 5.2, we have shown that transfer debiased estimators can enjoy sharper inference results if the contrast level $\delta_{1}^{(k)}$ is small. Now with other settings unchanged, we let $|\delta_{1}^{(k)}|=0.1$ and still compare the methods conducted in 5.2, where the results are shown in Table 2 and Figure 15.
From Table 2 and Figure 15, when $|\delta_{1}^{(k)}|$ is large, unfortunately our Debias_trans performs much worse than the target debiasing methods, which is in our expectation since Condition 7 is no longer satisfied. The density plots in Figure 15 show that the transfer debiased estimate is no longer normal, and the Shapiro-Wilk test rejects the normality hypothesis. In a similar way, the bootstrap approach in Algorithm 4 can also indicate us that debiasing on source studies will generate unreliable estimates. Therefore, we recommend only debiasing on the target study, namely the Debias_trans_tar approach, which can keep the coverage proportions with more stable interval lengths than Debias_dl.
8.2 Additional Simulation Results
We first introduce the simulation details for the motivating example in Section 1. The setting is similar to the simulation in [zhang2022transfer]. We set $p=500,s=10,n_{0}=200,\boldsymbol{\beta^{*}}=(\mathbf{0.5}_{s},\mathbf{0}_{p-s})^{\top},\boldsymbol{x}_{i}^{(0)}\sim\mathcal{N}(\mathbf{0}_{p},\boldsymbol{\Sigma}_{\boldsymbol{x}})$ with $\boldsymbol{\Sigma}_{\boldsymbol{x}}=(0.7^{|i-j|})_{1\leq i,j\leq p}$. We consider only one source study with sample size $n_{1}$ varying from 100 to 500, and we assume no parameter or covariate shift, i.e., $\boldsymbol{\delta}^{(1)}=0,\boldsymbol{x}_{i}^{(1)}\sim\mathcal{N}(\mathbf{0}_{p},\boldsymbol{\Sigma}_{\boldsymbol{x}})$. The key change lies in the residual distributions, where $\epsilon_{i}^{(0)}$ is standard normal, but $\epsilon_{i}^{(1)}$ follows one of the four types of distributions: 1) standard normal: $\mathcal{N}(0,1)$; 2) Cauchy distribution: $\mathcal{C}(0,3)$; 3) mixed Gaussian: $z\mathcal{N}(-3,0.5)+(1-z)\mathcal{N}(3,0.5)$, $z\sim\mathcal{B}ernoulli(\tau)$; 4) noisy normal: $\mathcal{N}(0,5^{2})$. It is worth noting that in this setting, the source study is included in the oracle transferable set by the previous definition [li2022transfer, tian2022transfer]. We compare four methods for estimating the target parameter $\boldsymbol{\beta^{*}}$ at the $0.2$-th quantile, which are: 1) L1-QR: $\ell_{1}$-penalized quantile regression [belloni2011l1] on the target data; 2) Pooling: $\ell_{1}$-penalized quantile regression on the entire dataset; 3) TransQR_2step: two-step transfer for quantile regression [huang2023estimation]; 4) TransSQR_2step: two-step transfer for smoothed quantile regression [zhang2022transfer]. We repeat these methods 50 times and report the average of $\ell_{2}$-estimation errors in Figure 1.
We conduct extensive experiments for parameter estimation and statistical inference under several circumstances. In particular, we consider the cases with strong covariate shift and $t$-distribution. Strong covariate shift means the covariates of each source study now follow $\mathcal{N}(\mathbf{0}_{p},\boldsymbol{\Sigma}_{\boldsymbol{x}}+\boldsymbol{A}_{\boldsymbol{x}}^{(k)\top}\boldsymbol{A}_{\boldsymbol{x}}^{(k)})$, where $\boldsymbol{A}_{\boldsymbol{x}}^{(k)}$ is a random matrix whose entries are $0.3$ with probability $0.3$ and 0 with probability $0.7$. $t$-distribution means we set $\epsilon_{i}^{(k)},k\in\{0\}\cup\mathcal{C}_{h_{2}}$ follow the shifted $t(3)$ distribution. Except the changes we mentioned, all other simulation settings are kept the same as those used in Figure 2 and 3 of the main manuscript.
Parameter estimation
  • Figure 5: heteroscedastic model, $K=5$, $h_{1}=5$.
  • Figure 6: homoscedastic model, $K=20$, $h_{1}=5$.
  • Figure 7: heteroscedastic model, $K=20$, $h_{1}=5$.
  • Figure 8: homoscedastic model, $K=5$, $h_{1}=15$.
  • Figure 9: heteroscedastic model, $K=5$, $h_{1}=15$.
  • Figure 10: homoscedastic model, $K=5$, $h_{1}=5$, strong covariate shift.
  • Figure 11: homoscedastic model, $K=5$, $h_{1}=5$, strong covariate shift, $t$-distribution.
  • Figure 12: heteroscedastic model, $K=5$, $h_{1}=5$, strong covariate shift, $t$-distribution.
  • Figure 13: homoscedastic model with varying $t_{1}$ values for TransQR. We set $\mathcal{C}_{h_{1}}=\{1,2\},\mathcal{C}_{h_{2}}=\{1,2,3,4,5\}$. We use five-fold cross-validation on the target study to compute the prediction error (quantile loss). The results are averaged over $20$ repetitions.
  • Figure 14: homoscedastic model with varying $t_{2}$ values for TransQR. We set $\mathcal{C}_{h_{1}}=\{1,2,3,4,5\},\mathcal{C}_{h_{2}}=\{1,2\}$. We use five-fold cross-validation on the target study to compute the prediction error (quantile loss). The results are averaged over $20$ repetitions.
Statistical inference
  • Figure 15 and Table 2: $|\delta_{1}^{(k)}|=0.1$.
  • Figure 16 and Table 3: strong covariate shift.
  • Figure 17 and Table 4: strong covariate shift.
Figure 5: Average $\ell_{2}$-error of various methods under heteroscedastic model with $K=5$.
Figure 6: Average $\ell_{2}$-error of various methods under homoscedastic model with $K=20$.
Figure 7: Average $\ell_{2}$-error of various methods under heteroscedastic model with $K=20$.
Figure 8: Average $\ell_{2}$-error of various methods under homoscedastic model with $h_{1}=15$.
Figure 9: Average $\ell_{2}$-error of various methods under heteroscedastic model with $h_{1}=15$.
Figure 10: Average $\ell_{2}$-error of various methods under homoscedastic model with strong covariate shift.
Figure 11: Average $\ell_{2}$-error of various methods under homoscedastic model and $t$-distribution with strong covariate shift.
Figure 12: Average $\ell_{2}$-error of various methods under heteroscedastic model and $t$-distribution with strong covariate shift.
Figure 13: Average prediction errors under different values of the threshold $t_{1}$ in TransQR.
Figure 14: Average prediction errors under different values of the threshold $t_{2}$ in TransQR.